OpenStudy (anonymous):

Can somone tell me how this works? ( √2 - √3 )² = A. 5 - 2√6 B. 5 - √6 C. 1 - 2√6 D. 1 - √2 E. 1 The answer is A. This is the explanation they give. Expand as for (a - b)2. (√2 - √3)(√2 - √3) = 2 - 2(√2 x √3) + 3 = 5 - 2 √6 I want to see how they came to the second and third step.

OpenStudy (anonymous):

by multiplication

OpenStudy (anonymous):

Multiplying what?

OpenStudy (anonymous):

\[(\sqrt{2}-\sqrt{3})(\sqrt{2}+\sqrt{3})\]

OpenStudy (anonymous):

you are multiplying each of the square roots in the first parenthesis by the ones in the second parenthesis. That makes sqrt 2 x sqrt 2+ sqrt 2 x -sqrt 3+ -sqrt 3 x sqrt 2+ -sqrt3 x -sqrt3

OpenStudy (anonymous):

sqrt 2xsqrt2 =2

OpenStudy (anonymous):

\[\sqrt{2}*\sqrt{2}-\sqrt{2}*\sqrt{3}-\sqrt{3}*\sqrt{2} -\sqrt{3}*\sqrt{3}\]

OpenStudy (anonymous):

there is positive with the last term 2 - 2(√2 x √3) + 3

OpenStudy (anonymous):

your last term is wrong its +sqrt3*sqrt3

OpenStudy (anonymous):

\[(a-b)^2=a^2+b^2-2ab \] so here put a=\[\sqrt{2} b= \sqrt{3}\] . u simplify urself and u'll get the answer.